OpenStudy (anonymous):

Im confused? n=15 L=(21/36)pi A. 7 B. 21 C.28 D. 62 E. 56 F. 14

OpenStudy (swissgirl):

I am confused too :) What's the topic?

OpenStudy (anonymous):

Circles: Circumferences and Arc Length

OpenStudy (swissgirl):

hmm Not sure what n stands for exactly

OpenStudy (anonymous):

me either

OpenStudy (swissgirl):

What does the question ask you to find?

OpenStudy (anonymous):

thats all it says is the equation :/

OpenStudy (anonymous):

do you have any idea what does the L stand for? Length?

OpenStudy (anonymous):

yes L is length

OpenStudy (anonymous):

Have you tried solving just for L? The result you get is:\[L=\frac{7 \pi}{180}\]Have you got any clue as to what n could be about?

OpenStudy (anonymous):

nope, i have no clue at all as of what i am doing

OpenStudy (anonymous):

lol.. is this the whole description of the problem at hand? Maybe there's more text to it, that could be very helpful..

OpenStudy (anonymous):

nope thats all i got, that why im confused

OpenStudy (anonymous):

Ok, there's one possibility and this is just hypothetical. It just came to me, I think that it could be asking for the angle in degrees. If that is the case, you would have to multiply this by\[\frac{180}{\pi}\]so when you multiply the result, L, by that factor you get,\[\frac{7 \pi}{180} \times \frac{180}{\pi} = 7\]This could mean the answer is A. But I can't assure you, because, well arc-lengths are not in degrees. Have you been doing some like this in class? The conversion and stuff..